# quality of service 5G fifth generation

### v2

@@ V-A Initial points and performance of training the model
Figure 4 shows the training and validation loss values. The decreasing trend of loss values indicates that the SS-VAE method could accurately estimate the target values (power of transmission, association indicators). The model can map the input channel gain matrix to the power of transmission and the indicators by employing the encoder part of the VAE’s to learn the underlying information.
The performance metrics for the test data, as given in Table V, clearly indicate that SS-VAE method obtained optimal performance in the test dataset, demonstrating its generalization ability. The $R^{2}_{Score}$ value for the test data indicates that our model successfully explains a significant portion of the variance in estimating the power of transmissions. This improvement can be attributed to the utilization of VAE, which has been proven to enhance generalization in various related tasks. Furthermore, the pre-training method involving contrastive loss has improved our model’s ability to learn the underlying features, reducing generalization errors.
TABLE V: Performance metrics of the SS-VAE model
Performance Metric MAE $R^{2}_{Score}$ Cosine Similarity
@@
V-B Comparison of the algorithms
We evaluated the performance of our SS-VAE method compared to the ESA algorithm, focusing on how the number of labeled samples used for training affects the results, as shown in Figure 5. We analyze the aggregated throughput under varying numbers of labeled data points obtained by SS-VAE algorithm and the ESA. As depicted in Figure 5a, SS-VAE achieves a similar aggregated throughput compared to the ESA. Moreover, the discrepancy between SS-VAE and the ESA decreases significantly as the number of labeled data points increases. The difference in aggregated throughput between these methods is $2.1$ for $2000$ labeled samples, which will be decreased using more labeled samples.
(a) Agg. throughput vs. Number of labeled data.
(b) Association Error vs. Number of labeled data.
Figure 5: Comparison of the algorithms (SS-VAE vs. ESA).
We also evaluate the performance of the SS-VAE and ESA based on binary variables related to user association ($\alpha_{u,b,s}$) and PRB allocation ($\beta_{u,b,m,s}$). Figure 5b displays the ratio of the absolute sum of differences in these binary variables between the SS-VAE and ESA to the total number of them ($\frac{|\alpha_{ESA}-\alpha_{SS-VAE}|+|\beta_{ESA}-\beta_{SS-VAE}|}{\text{\# Number of }(\alpha\&\beta)}$), as the number of labeled samples increases. The difference between them reaches its minimum at 2000 samples (4.7% error) and is expected to decrease significantly as more labeled samples are used.
@@ V-B Comparison of the algorithms
V-C Performance results
(a) Aggr. throughput vs. number of UEs.
(b) Aggr. throughput vs. number of UEs ($P_{O-RU}$).
Figure 6: Aggregated throughput of the SS-VAE vs. Number of UEs.
@@ V-C Performance results
Figure 6 illustrates the aggregate weighted rate across different numbers of UEs as achieved by three algorithms: SS-VAE, ESA, and DQN. Figure 6a demonstrates the aggregated throughput obtained by SS-VAE, ESA, and DQN for different number of UEs. The difference in performance between SS-VAE and ESA is consistent across all UEs. Moreover, SS-VAE outperforms DQN in aggregated throughput across all UE. In addition, the aggregated throughput increases initially as more UEs join the service for all three examined algorithms. However, it reaches a plateau after 30-35 UEs per service due to power limitations and interference constraints. Figure 6b shows the aggregated throughput achieved by the SS-VAE algorithm at different maximum power levels of O-RU and different numbers of UEs. In resonation with [15], increasing the maximum power of O-RU and the number of UEs leads to higher aggregated throughput. The highest aggregated throughput is observed when the maximum power of O-RU is at its highest.
(a) Aggr. throughput vs. O-RU power.
(b) Aggr. throughput vs. Power of slice.
Figure 7: Aggregated throughput of SS-VAE vs. Power values.
@@
Figure 7 displays the aggregated throughput for two specified service types (eMBB and URLLC), which has been obtained using the SS-VAE and DQN algorithms, about the maximum power of slice and RU. Figure 7a and Figure 7b both illustrate the aggregate weighted throughput achieved by SS-VAE and DQN under differetn maximum power levels for RU and slices, respectively, in eMBB and URLLC services. Both figures demonstrate a direct correlation between increased maximum power levels and higher weighted throughput. Across all scenarios, the SS-VAE algorithm consistently outperforms the DQN algorithm, providing higher aggregate throughput for both service types, regardless of whether the power levels pertain to the RU or the individual slices. Due to their need for higher data rates, eMBB service type exhibits higher aggregate throughput than URLLC services. This consistent performance advantage underscores the effectiveness of the SS-VAE algorithm in optimizing resource allocation across different power settings.
@@
Although ESA ensures a globally optimal solution, its computational complexity makes it impractical for real-world applications. DQN, while powerful, faces challenges related to convergence time, exploration efficiency, and sample requirements. Our SS-VAE method employs the strengths of supervised and unsupervised learning to provide a scalable and efficient solution for resource allocation in O-RAN, outperforming these benchmarks in real-world scenarios. This contrasts to DQN, which may struggle with generalization due to its reliance on extensive exploration and sample efficiency. Additionally, DQN often requires significant training time to converge to an optimal solution, especially in environments with high variability and complex constraints. It may also face challenges with exploration efficiency, often needing many iterations to fully explore the solution space, resulting in suboptimal performance in dynamic conditions.
VI COMPUTATIONAL COMPLEXITY ANALYSIS
@@ VI-1 ESA
Total $\text{Total Cost}=(|U|\times|B|)\times(M+1)!\times P_{\text{levels}}$
Where $(M+1)!$ represents the worst case of assigning UE to PRB or not, ESA needs to scale better with increasing network size and complexity due to its exponential growth in computational requirements. Despite its guarantee of finding the global optimal solution, it is impractical for real-time applications and large-scale networks.
VI-2 DQN
@@ VI-4 Comparison and Implications
VII CONCLUSION
This paper presents a novel model for the downlink of the O-RAN system, designed to coordinate multiple independent xAPPs to address the resource allocation problem with network slicing in the O-RAN architecture. It focuses on optimizing the aggregate weighted throughput by managing RU associations, power allocation, and PRBs to support two central 5G services, eMBB, and URLLC.
Due to the non-convex and NP-hard nature of the resource allocation optimization problem, we proposed a semi-supervised learning approach to solve the resource allocation problem efficiently. Our algorithm begins by reformulating the problem as an optimization problem. The training phase involves supervised learning using a VAE to estimate power transmission and association indicators, followed by unsupervised learning using a contrastive loss approach to improve the model’s generalization and robustness.
We compared the SS-VAE method with benchmark algorithms ESA and DQN in various scenarios and discussed their computational complexity. The results indicate that our method performs similarly to ESA while outperforming DQN. We also examined scenarios for eMBB and URLLC services, considering their specific QoS requirements. Through simulations, we demonstrated the efficiency of our model, making it a robust candidate for resource allocation in O-RAN systems. Given the dynamic and complex nature of O-RAN environments, the sample efficiency of our model allows it to adapt to changing conditions without extensive retraining, ensuring high performance with fewer labeled data samples and enabling faster deployment and response times. The SS-VAE model provides robustness, reliability, and generalization from limited data.
In our future research, we aim to explore several areas. Firstly, we intend to evaluate the end-to-end delay experienced by individual UEs and study the impact of deployment costs associated with network function placement. Addressing UE mobility is crucial in practical scenarios. Furthermore, we will explore alternative ML and optimization algorithms to improve our resource allocation approach, paving the way for new research opportunities in this dynamic field.

### v3

@@ 6.2 Performance of Model Training
This section describes the performance of the trained model. Figure 5 shows the training and validation loss values. The decreasing trend of loss values indicates that the SS-VAE method could accurately estimate the target values (power of transmission, association indicators). The model can map the input channel gain matrix to the power of transmission and the indicators by employing the encoder part of the VAE to learn the underlying information.
The performance metrics for the test data, as given in Table VI, clearly indicate that the SS-VAE method obtained optimal performance in the test dataset, demonstrating its generalization ability. The $R^{2}_{Score}$ value for the test data indicates that our model successfully explains a significant portion of the variance in estimating the power of transmissions. This improvement can be attributed to the utilization of VAE, which has been proven to enhance generalization in various related tasks. Furthermore, the pre-training method involving contrastive loss has improved our model’s ability to learn the underlying features, reducing generalization errors.
TABLE VI: Performance metrics of the SS-VAE model
Performance Metric MAE $R^{2}_{\text{Score}}$ Cosine Similarity
@@
6.3 Comparison of the Algorithms
(a) Aggregated throughput of SS-VAE and ESA.
(b) Association error of SS-VAE and ESA.
(c) Aggregated throughput vs. number of UEs.
(d) Aggregated throughput vs. number of UEs ($P_{O\text{-}RU}$).
(e) Aggregated throughput vs. O-RU power.
(f) Aggregated throughput vs. Power of slice.
Figure 6: Simulation results showing various performance metrics and algorithm comparisons for SS-VAE, ESA, and DQN models.
We evaluated the performance of our SS-VAE method compared to the ESA algorithm, focusing on how the number of labeled samples used for training affects the results, as shown in Figure 6a. We analyze the aggregated throughput under varying numbers of labeled data points obtained by the SS-VAE algorithm and the ESA. As depicted in Figure 6a, SS-VAE achieves a similar aggregated throughput compared to the ESA. Moreover, the discrepancy between SS-VAE and the ESA decreases significantly as the number of labeled data points increases. The difference in aggregated throughput between these methods is $2.1$ for $2000$ labeled samples, which will be decreased using more labeled samples.
We also evaluate the performance of the SS-VAE and ESA based on binary variables related to user association ($\alpha_{u_{s}}^{b}$) and PRB allocation ($\beta_{u_{s},m}^{b}$). Figure 6b displays the ratio of the absolute sum of differences in these binary variables between the SS-VAE and ESA to the total number of them $(|\alpha_{ESA}-\alpha_{SS-VAE}|+|\beta_{ESA}-\beta_{SS-VAE}|)/\text{\# Number of }(\alpha\&\beta)$, as the number of labeled samples increases. The difference between them reaches its minimum at 2000 samples (4.7% error) and is expected to decrease significantly as more labeled samples are used.
@@ 6.3 Comparison of the Algorithms
6.4 Performance Results
@@
Figure 6c demonstrates the aggregated throughput obtained by SS-VAE, ESA, and DQN for different number of UEs. The difference in performance between SS-VAE and ESA is consistent across all UEs. Moreover, ESA achieves higher aggregate throughput than SS-VAE, outperforming DQN in total throughput across all UEs. In addition, the aggregated throughput increases initially as more UEs join the service for all three examined algorithms. However, it reaches a plateau after 30-35 UEs per service due to power limitations and interference constraints. Figure 6d shows the aggregated throughput achieved by the SS-VAE algorithm at different maximum power levels of O-RU and different numbers of UEs. In resonation with [17], increasing the maximum power of O-RU and the number of UEs leads to higher aggregated throughput. The highest aggregated throughput is observed when the maximum power of O-RU is at its highest.
@@
Figures 6e and 6f show the aggregated throughput for two service types—eMBB and URLLC—obtained using the SS-VAE and DQN algorithms. Specifically, Figure 6e illustrates the weighted aggregate throughput as a function of the maximum power of the RU, while Figure 6f presents the throughput variation concerning the maximum power allocated per slice. Both figures compare the performance of SS-VAE and DQN across different power constraints for these services. They demonstrate a direct correlation between increased maximum power levels and higher weighted throughput. Across all scenarios, the SS-VAE algorithm consistently outperforms the DQN algorithm, providing higher aggregate throughput for both service types, regardless of whether the power levels pertain to the RU or the individual slices. Due to their need for higher data rates, eMBB service type exhibits higher aggregate throughput than URLLC services. This consistent performance advantage underscores the effectiveness of the SS-VAE algorithm in optimizing resource allocation across different power settings.
@@
Although ESA ensures a globally optimal solution, its computational complexity makes it impractical for real-world applications. DQN, while powerful, faces challenges related to convergence time, exploration efficiency, and sample requirements. Our SS-VAE method employs the strengths of supervised and unsupervised learning to provide a scalable and efficient solution for resource allocation in O-RAN, outperforming these benchmarks in real-world scenarios. This contrasts with DQN, which may struggle with generalization due to its reliance on extensive exploration and sample efficiency. Additionally, DQN often requires significant training time to converge to an optimal solution, especially in environments with high variability and complex constraints. It may also face challenges with exploration efficiency, often requiring many iterations to fully explore the solution space, which can result in suboptimal performance under dynamic conditions.
7 COMPUTATIONAL COMPLEXITY ANALYSIS
@@ 7.0.1 ESA
Total $\text{Total Cost}=(|B|^{|U|})\times\prod_{b}(\sum_{s}\sum_{u}\alpha_{u_{s}}^{b}+1)^{M}\times P_{\text{levels}}$
Where $\prod_{b}(\sum_{s}\sum_{u}\alpha_{u_{s}}^{b}+1)^{M}$ represents the worst case number of possible PRB assignments to UEs without intra-cell interference in each RU $b$, ESA needs to scale better with increasing network size and complexity due to its exponential growth in computational requirements. Despite its guarantee of finding the global optimal solution, it is impractical for real-time applications and large-scale networks.
7.0.2 DQN
@@ 7.0.4 Comparison and Implications
8 Conclusions
This paper introduced SS-VAE, a novel GAI framework for optimizing downlink O-RAN resource allocation and network slicing. SS-VAE coordinates xApps to manage RU associations, power, and PRBs, maximizing weighted throughput for eMBB and URLLC services. To tackle the NP-hard optimization challenge, SS-VAE employs a unified semi-supervised DNN. This architecture synergistically integrates a VAE—whose generative capabilities are honed using ESA-derived optimal parameters to determine allocation decisions—with a contrastive loss function that enhances representation learning, generalization, and robustness using both labeled and unlabeled data.
Evaluations against the ESA and DQN benchmarks demonstrated SS-VAE’s ability to achieve near-optimal performance, comparable to ESA and significantly outperforming DQN while incurring the reduced computational cost. Our simulations confirmed the efficacy of SS-VAE in managing eMBB/URLLC QoS across diverse O-RAN scenarios. As a semi-supervised GAI framework, SS-VAE offers notable sample efficiency, robust generalization from limited data, and adaptability to dynamic conditions, marking it as a compelling solution for practical O-RAN deployments.
Future research will address end-to-end UE delay, investigate deployment costs, and incorporate UE mobility. We also aim to explore enhanced GAI techniques and other advanced AI or hybrid optimization algorithms to improve further resource allocation strategies in dynamic O-RAN systems, thereby opening new avenues for research in this evolving field.